OpenStudy (anonymous):

-Calculate [H3O+] for the 0.10 M CH3COOH solution which has a pH = 2.88. -What percent of the weak acid is dissociated? *Can someone just help me figure out how i would approach the problem?? Thanks!(:

OpenStudy (aaronq):

well you know \([H_3O^+]\) from pH, so just find what percent \([H_3O^+]\) is from\([CH_3COOH]\) (which is 0.1 M).

OpenStudy (anonymous):

Thanks for replying!! ok so I got 1.32x10^-3 for the [H3O+] .. How would I find what percent of that is from CH3COOH?

OpenStudy (aaronq):

\(\dfrac{[H_3O^+]}{[CH_3COOH]}*100\%\)

OpenStudy (aaronq):

hm no wait. that seems wrong

OpenStudy (anonymous):

so for the bottom part, are you referring to the molecular mass in moles?

OpenStudy (anonymous):

i was thinking.. i don't actually remember seeing that formula in class, but you're the expert! haha(:

OpenStudy (aaronq):

hm it's not really a formula, it's like if you have 10 cats and you let 2 lose, what percent of cats did you let lose? \(\dfrac{2}{10}*100\%\)=20% same kind of thing

OpenStudy (aaronq):

i think it works: 0.0013182567385564/0.1*100%=1.3182567385564%

OpenStudy (anonymous):

ohh ok! so that's for the second part though right? and the first part i put in what i got for h30 right?

OpenStudy (anonymous):

Wait nvm, I got it! hahaha.. ok so let me just put in the second part and see if that's right!

OpenStudy (anonymous):

Totallly worked, thank you sooooooooooo soooooo much! Hahaha.. Seriously, most amazing person on here! Thanks!!(:

OpenStudy (aaronq):

okay awesome ! no problem at all ! i wasn't sure if that was right for a min lol

OpenStudy (anonymous):

Oh come on now, you're always right! Hahaha, well once again, thanks a lot!!(: